OpenStudy (anonymous):

a grocer wants to mix two kinds of nuts. one kind sells for $0.95 per pound, and the other sells for $2.55 per pound. He wants to mix a total of 22 pounds and sell it for $1.20 per pound. How many pounds of each kind should he use in the new mix?

OpenStudy (anonymous):

Let's call the first nut, nut A. Nut A sells for .95 per pound. The second nut we'll call nut B, and it sells for 2.55 per pound. Together he wants 22 pounds, so A + B = 22. He wants them together to be worth 1.20 per pound, which means that the cost of all the nuts divided by the weight should be 1.20, so (.95A + 2.55B)/22 = 1.20

OpenStudy (anonymous):

That gives you two equations in two unknowns. A+B =22 (.95/22)A + (2.55/22)B = 1.20 Now just solve one equation for one of the variables, substitute that into the other equation and solve for the variable.

OpenStudy (anonymous):

Then take that value, substitute it back into one of the original equations and solve for the remaining variable. Kablam, sucka.